# Closed Loop Interactive Embodied Reasoning for Robot Manipulation

### v1

@@ Abstract
Embodied reasoning systems integrate robotic hardware and cognitive processes to perform complex tasks typically in response to a natural language query about a specific physical environment. This usually involves changing the belief about the scene or physically interacting and changing the scene (e.g. Sort the objects from lightest to heaviest). In order to facilitate the development of such systems we introduce a new simulating environment that makes use of MuJoCo physics engine and high-quality renderer Blender to provide realistic visual observations that are also accurate to the physical state of the scene. Together with the simulator we propose a new benchmark composed of 10 classes of multi-step reasoning scenarios that require simultaneous visual and physical measurements. Finally, we develop a new modular Closed Loop Interactive Reasoning (CLIER) approach that takes into account the measurements of non-visual object properties, changes in the scene caused by external disturbances as well as uncertain outcomes of robotic actions. We extensively evaluate our reasoning approach in simulation and in the real world manipulation tasks with a success rate above 76% and 64%, respectively.
@@ I Introduction
The research efforts in developing systems capable of high-level perception and reasoning [1, 2, 3, 4, 5, 6] have been increasing in recent years. Many robotics systems include an agent that is required to perform a task with a specified goal given visual observations or instructions in natural language [7, 8, 1, 4, 9]. Such tasks requires a simulation environment for generating data, or on-line agent training, together with an evaluation process. Existing environments [10, 11] differ in terms of physics simulation engines and the quality of visual observations. High computational load of realistic rendering forces a compromise between real-time physics calculation and visual quality of the simulation. Realistic synthetic visual data are crucial for development of robotics systems but the collection of such data is often restricted by a cumbersome setup process and real time robot operations [12]. More powerful simulators, datasets and benchmarks are needed for a closed-loop setup where changes to the world or physical measurements (e.g., weight, elasticity, occlusions, etc.) should be accounted for as directly affecting the next step in action planning.
Figure 1: An example task presenting capabilities of the proposed environment: synthetic scene generation, instruction generation, execution of symbolic actions for manipulation followed by physics calculation and realistic rendering. Symbolic actions with corresponding targets marked in the image. Best viewed zoomed.
In this paper, we emphasise the following contributions [^1] The environment, the benchmark and the reasoning approach are released on the project webpage https://michaal94.github.io/CLIER:
$\circ$ We introduce MuBlE, a novel modular environment for simulating and executing interactive manipulation tasks characterised by high-quality visual rendering and accurate physics calculation. It employs MuJoCo for physics and Blender for rendering, which we refer to with MuBlE (MuJoCo+Blender Environment). It can generate multimodal data based on scene specification and instruction templates. An example scene and task that can be executed in this environment is shown in Fig. 1. We believe to be the first work focusing on task planning in robotics manipulation while preserving accurate physics modelling between the bodies, in contrast to discrete action spaces such as in ManipulaThor [8] and ALFRED [4].
$\circ$ We propose SHOP-VRB2, a new benchmark for evaluating embodied closed loop reasoning systems for robotic manipulation. It includes 10 types of benchmarking tasks for single- and multi-step tabletop manipulations that require combined reasoning between visual observations (recognising attributes of objects and their relations) and physical measurements (acquiring properties obtainable through interaction only, such as weight or stiffness). SHOP-VRB2 extends [13] with a challenging combination of various modalities (language, vision, and manipulation) for benchmarking Visual Reasoning, VQA (e.g. [14]) (language + vision), and Embodied question answering (EQA) (e.g. [6, 15]) (language + vision + navigation) approaches.
$\circ$ We propose a closed-loop neuro-symbolic interactive reasoning approach (CLIER), that can plan, evaluate and execute a given task (e.g., Place the heaviest mug on the box). The highlight of our approach is that it can plan for perception of invisible object properties (e.g., weight or stiffness) and account for new measurements to adjust the plan. Our reasoning agent is able to recover from a wide range of manipulation failures caused by external disturbances and uncertain outcomes of manipulation actions.
@@
$\circ$ We evaluate our proposed Closed Loop Interactive Embodied Reasoning (CLIER) within MuBlE using the proposed SHOP-VRB2 synthetic benchmark and validate it via comparative experiments between simulated and real-world YCB [16] objects.
In the remainder of this paper, we first review the existing simulators and justify the need for a new environment. We then present the proposed MuBlE and CLIER in Sec. III. Sec. IV introduces our SHOP-VRB2 benchmark. Results and discussion are presented in Sec. V.
@@ II Related work
TABLE I: Comparison of the MuBlE’s capabilities with the related simulators for the robotic experiments. Our MuBlE can be characterised by high quality rendering with Blender, accurate physics with MuJoCo, included visible and non-visible states, various action spaces available, targeting tabletop manipulation, modular and easily extendable.
Simulator Rendering Physics Motion planner Non-kin. states Non-visible states Action space Speed Scale Open-source
library material quality library supports
AI2-THOR [10] Unity textures ++ Unity rigid dyn./animation $\checkmark$ Discrete + room $\checkmark$
ManipulaTHOR [8] Unity textures ++ Unity AI2THOR+manip. Discrete + room $\checkmark$
ThreeDWorld [17] Unity/V-Ray (off.) Configurable $+$ Unity+FLEX rigid/particle dyn. Continuous ++ house
Robosuite [18] MuJoCo/NVISII Configurable ++ MuJoCo rigid/artic. dyn. $\checkmark$ Continuous ++ tabletop $\checkmark$
ISAACSim [11] Omniverse RTX Configurable ++ PhysX rigid/artic. dyn. $\checkmark$ (Disc.)&Cont. ++ house
Habitat 2.0 [9] Magnum 3D scans/PBR + PyBullet rigid/artic. dyn. Continuous $+$ house $\checkmark$
CoppeliaSim [19] OpenGL Gouraoud shad. + PyBullet rigid/artic. dyn. $\checkmark$ (Disc.)&Cont. ++ room
iGibson [20] PyRender/OpenGL PBR shad. + PyBullet rigid/artic. dyn. $\checkmark$ $\checkmark$ Continuous ++ house $\checkmark$
MuBle (ours) MuJoCo/Blender procedural $+$ MuJoCo rigid/artic. dyn. $\checkmark$ $\checkmark$ $\checkmark$ Disc.& Cont. + tabletop $\checkmark$
Simulation environments for manipulation focused on combined visual and language robotic tasks. State-of-the-art simulators ManipulaTHOR [8] (extension of AI2-THOR [10]), CoppeliaSim [19], or iGibson [20] provide non-photorealistic rendering, lacking accurate shades and reflections. This is in contrast to Blender that can generate photorealistic scenes with procedural object models. ISAACSim [11] is a resource-intensive, closed-source simulator with limited customisation. ThreeDWorld [17] also makes use of a commercial renderer and does not provide the infrastructure for robotics manipulation tasks, including motion planners. Unlike prior works, we pay particular attention to conditional action planning and execution in robotics manipulation while preserving accurate physics modelling between the bodies, and offer both continuous and discrete action spaces, in contrast to discrete action spaces such as in ThreeDWorld [17] and ManipulaThor [8], or pure continuous control of the robot as seen in Robosuite [18]. For example, ManipulaThor [8] considers grasping as a discrete action, while we enable to execute sequences of various custom primitive actions (e.g., approach, gripper closing, lifting, etc.) or continuous control of the robot. From these simulators, only iGibson offers non-kinematic continuous and discrete object states e.g. temperature or burned. Our MuBlE environment differs from the aforementioned manipulator simulators by a unique and open-source combination of 1) high-quality visual data and realistic physics modelling, which is a key advantage when attempting sim2real transfers (as shown in our experiments), 2) ability to select between operating on continuous action space or usage of primitive actions, and 3) considering both visible and non-visible continuous object states (e.g., weight, or stiffness). We present a comparison between different simulators in Table I.
Benchmarks for language-conditioned manipulation. The advantages of MuBlE environment enabled us to prepare a unique benchmark for closed-loop embodied reasoning (SHOP-VRB2) that extends SHOP-VRB [13]. SHOP-VRB is a dataset closely related to the robotic manipulation scenarios considered in this work. It features various objects suitable for robotic manipulation that can be easily obtained. There are many high quality 3D models of various instances of these objects which make them suitable for generating synthetic scenes via Blender rendering that includes procedurally generated materials rather than simply texturing the mesh. We therefore extend it to a dataset and a benchmark SHOP-VRB2 that enables development and testing of systems capable of reasoning on non visual attributes and performing interactive tasks.Compared to other benchmarks, such as ALFRED [4] (based on RoboTHOR), a benchmark for understanding natural language instructions in robotics, VLMbench [21], based on RLbench [22] and CoppeliaSim [19], that adds linguistic commands and realises automatic complex tasks builders, IKEA Furniture Assembly [23], that provides a benchmark for various assembly tasks and CALVIN benchmark [24] offering natural language instructions for long-horizon manipulation tasks, our SHOP-VRB2 benchmark requires long-horizon closed-loop embodied reasoning, that involves manipulation of the world to acquire knowledge about non-visual object properties in order to complete the task or answer a query.
@@
Symbolic approaches gained interest in visual reasoning tasks and several deep learning systems incorporate neuro-symbolic programs. CLEVR-IEP [25] suggested predicting symbolic programs to be executed on disentangled scene representation, and was followed by NS-VQA [26], and NS-CL [27]. V2A [28] introduced a symbolic approach for robotic tasks using only the initial state of the scene, and was evaluated in a real robotic environment [6], however, only open-loop single-step tasks were demonstrated. Symbolic approaches were also considered in Task and Motion Planning systems (TAMP) that were often based on the formal language PDDL [29]. Classic TAMP approaches rely on predefined rules and known dynamic models [30, 31, 32] but learning based methods are increasingly replacing handcrafted heuristics in TAMP methods [33, 34, 35]. PDDLStream [36] extends PDDL to add any sub-symbolic model in a black-box way, which also opens it to the use of learning-based methods. However, it works only on a symbolic level and does not close the loop with the real world. Deep Visual Reasoning [37] directly predicts task plans by considering initial observation only. Regression Planning Networks (RPN) [38] propose a task planning model which performs regression in symbolic space using neural networks. RPN are expanded in [39] by utilising regression planning on scene graph representations to estimate the next subgoal based on the current scene and a given scene graph, i.e. solving only subpart of the overall task that our method deals with. The interaction with the real world that follows a linguistic query (e.g.measure the weight of the mug, find the sofa) is explored in embodied question answering approaches [15, 6, 40]. In [6], language-conditioned visual reasoning is combined with manipulation. Navigation of agents based on vision and language was investigated in VLN [1]. However, both approaches considered tasks that do not react to incoming measurements, i.e. reasoning is not conditioned on new observations.
Figure 2: A diagram of CLIER reasoning and MuBlE environment including interaction between the two, and use of SHOP-VRB2 benchmark. Symbols for transferred data: $\mathtt{T}$ text of the query, $\mathtt{\tilde{G}}$ prediction of scene graph elements, $\mathtt{G}$ current scene graph, $\mathtt{S}$ subgoal (symbolic program requiring physical measurements), $\mathtt{I}$ image, $\mathtt{P}$ physical observations, $\mathtt{C}$ control signal, $\mathtt{A}$ primitive action to take, $\mathtt{R}$ returned result.
Recently, several works utilised large-langauge models (LLMs) to generate action plans for robot manipulation tasks (eg.,  [41], [42]). While they show impressive versatility and adaptability to novel environments, they suffer from lack of explainability, verifiability, and repeatability. LLMs require careful prompt engineering and filtering of outputs before the robot can execute them. Furthermore, such approaches do not capture the task and trajectory details and do not enable fast-loop reactiveness. These problems have to be resolved to broadly deploy LLMs for embodied reasoning. Using data created with MuBlE environment, we train and evaluate a novel CLIER reasoning agent that can act on a natural language query in a real scene. Moreover, we present closed-loop reasoning experiments in simulation and real tabletop scenes. A highlight of our approach is that it plans for perception of invisible object properties, e.g. when stacking objects by weight. Moreover, unlike classical methods, our reasoning agent is executed on every key frame, which allows it to recover from a wide range of manipulation failures (e.g., unsuccessful grasping, or falling object during manipulation).
III MuBlE Environment and CLIER Reasoning
@@
In this section, we present our MuBlE environment for simulating manipulation tasks and reasoning CLIER approach that makes use of this environment. The main modules and data flow are presented in Fig. 2.
MuBlE is built on $\mathtt{robosuite}$ [18] which is a simulation framework suitable for creating robotic environments inside the physics engine MuJoCo. We equip our environment with high quality rendering powered by Blender. It is designed for a generic tabletop scenario with a single robotic manipulator and a gripper. We use the same set of robots and grippers as $\mathtt{robosuite}$. We provide a set of object models in MuJoCo and their counterparts in Blender along with templates to create new items.
We first introduce the scene graph in Sec. III-A and action planning in Sec. III-B that is a part of CLIER. We then present two main components of the environment and reasoning approach: action (render) loop in Sec. III-C, and physics loop in Sec. III-D. Sec. III-E summarises our reasoning pipeline.
@@
Figure 3: (Left) Examples of visual observations (selected frames) generated for the actions corresponding to the instruction: Stack the lightest of metal objects on the yellow object. Note that metal cans were initially picked up to measure their weight. Further, according to the measurement taken, the heavier of the cans was put down and the lighter one was picked up and stacked on the yellow plate. (Right) Example scenes and corresponding instructions in natural language generated with MuBlE (in the dataset, instructions left to right belong to tasks 7, 3, and 1 in Tab. II).
III-A Scene Graph
@@
A real scene can be captured by a camera or a synthetic one can be rendered by Blender based on a detailed scene description. Scene graph $\mathtt{G}$ is generated for each key frame $\mathtt{I}$ starting from the first image of the scene. A keyframe is rendered after a motion corresponding to the primitive action is finished e.g. moving from object 1 to object 2, approaching grasping pose, closing the gripper, etc. We have chosen the keyframe based approach as it allows us to use Blender for rendering frames, which have sufficient quality for sim2real transfer, and sufficient speed to allow for recovering from failures with the closed loop approach.
@@
To parse the scene into a set of objects with attributes $\mathtt{\tilde{G}}$ we apply an instance segmentation with Mask R-CNN [43], followed by classification of object attributes with ResNet [44] and pose regression with CosyPose [45]. Scene graph is a sequence of feature vectors corresponding to each object in the scene including whether the object is in the gripper or whether the gripper is closed. We use simple geometrical heuristics to create or update the scene graph $\mathtt{G}$ for keyframe $\mathtt{I}$, similarly to [39].
High-quality rendering of every frame can be used for training but it is too slow for real-time experiments. However, many tasks are achievable with low rate visual reasoning and fast feedback loops from other sources of information (e.g. force control). Example key frames rendered during execution of a task are presented in Fig. 3 (Left).
III-B Symbolic program generator and action planner
As the reasoning method we propose a multi-staged approach inspired by Neural Symbolic VQA approaches [13, 26]. Symbolic program generator is implemented as a Seq2Seq network [46] that given translates the natural language query $\mathtt{T}$ into a sequence of symbolic programs that are evaluated on the scene graph $\mathtt{G}$ for keyframe $\mathtt{I}$, producing a subgoal $\mathtt{S}$ for the task e.g. measuring weight. We employ a transformer as the action planner to predict the next primitive action $\mathtt{A}$ and its target, given the current scene graph $\mathtt{G}$ and subgoal $\mathtt{S}$.

### v2

@@ Abstract
Embodied reasoning systems integrate robotic hardware and cognitive processes to perform complex tasks, typically in response to a natural language query about a specific physical environment. This usually involves changing the belief about the scene or physically interacting and changing the scene (e.g. sort the objects from lightest to heaviest). In order to facilitate the development of such systems we introduce a new modular Closed Loop Interactive Embodied Reasoning (CLIER) approach that takes into account the measurements of non-visual object properties, changes in the scene caused by external disturbances as well as uncertain outcomes of robotic actions. CLIER performs multi-modal reasoning and action planning and generates a sequence of primitive actions that can be executed by a robot manipulator. Our method operates in a closed loop, responding to changes in the environment.
Our approach is developed with the use of MuBle simulation environment and tested in 10 interactive benchmark scenarios. We extensively evaluate our reasoning approach in simulation and in real-world manipulation tasks with a success rate above 76% and 64%, respectively.
@@ I Introduction
The research efforts in developing systems capable of high-level perception and reasoning have increased in recent years ([1, 2, 3, 4, 5, 6, 7, 8, 9]), including models that utilize large pretrained vision-language models for action planning such as [10, 11, 12]. However, the ability of these models to reason iteratively through complex problems is restricted by the necessity to ground the reasoning about robot policies in sensory observations and the robot state. Even if some of the above-mentioned reasoning agents are able to generate and execute a plan, either their execution, their planning, or both are open loop (e.g. [6, 8]). This means that they do not account for failures during the execution (open-loop execution) and cannot react to the knowledge acquired during execution, thus cannot solve long-term planning that requires to collect and maintain a set of sequential observations that change the action plan such as ’find the softest mattress’ (open-loop planning). Furthermore, they typically reason only over visual observations and goals, which do not require physical measurements via interaction with the environment (see [9], where exploratory actions to find physical object properties are selected using inference in a Bayesian network). Such methods are unable to robustly execute tasks such as passing the lightest bottle or finding the object with the coarsest surface. Although some recent models (e.g., [7], [13], [5]) incorporate closed-loop reasoning, they focus on short-term tasks, e.g., manipulating an object. Robust and reactive long-horizon embodied reasoning remains an open problem.
Figure 1: A diagram of CLIER reasoning implemented in MuBlE environment including interaction between the two and use of SHOP-VRB2 benchmark. Transferred data: $\mathtt{T}$ – text of the query, $\mathtt{\tilde{G}}$ – prediction of scene graph elements, $\mathtt{G}$ – current scene graph, $\mathtt{S}$ – subgoal (symbolic program requiring physical measurements), $\mathtt{I}$ – image, $\mathtt{P}$ – physical observations, $\mathtt{C}$ – control signal, $\mathtt{A}$ – primitive action to take, $\mathtt{R}$ – returned result.
The main contribution of this article is the proposed Closed-Loop neuro-symbolic Interactive Reasoning framework (CLIER) [^1] The code and its documentation are released on the project webpage https://michaal94.github.io/CLIER. that can plan, evaluate, and execute a given task (e.g., Move the heaviest metal object to the right) (see Fig. 1). The highlight of our neuro-symbolic approach is that it can plan for perception of physical object properties (e.g., weight or stiffness) and account for new measurements to adjust the plan. CLIER processes visual inputs to identify objects and their poses, updating a scene graph before passing the information to an action planner. The planner selects the next robotic action, e.g., move or weigh (see Sec.III-E). Each action runs in a fast feedback loop (see Primitive Action Controller in Fig.1). We refer to all actions as primitive actions and any resulting observations—potentially from physical interactions—as measurements. Our reasoning agent is able to recover from a wide range of manipulation failures.
Figure 2: An example long-horizon manipulation task from SHOP-VRB2 implemented within the MuBlE environment [14]. A synthetic scene is rendered in Blender every keyframe, followed by the execution of the symbolic manipulation action planned by CLIER.
@@
We evaluate our proposed Closed Loop Interactive Embodied Reasoning (CLIER) within the MuBlE simulation environment [14] using the SHOP-VRB2 synthetic benchmark (see an example task in Fig. 3) and validate it via comparative experiments between simulated and real-world YCB-VRB benchmark. Both of these datasets require combined reasoning between visual observations (recognising attributes of objects and their relations) and physical measurements (acquiring properties obtainable through interaction only, such as weight or stiffness).
@@ II Related work
Symbolic approaches to robot planning date back to the STRIPS planner [15], but there has been renewed interest in the possibility of incorporating unprocessed sensory, in particular visual, information. Several deep learning systems incorporate neuro-symbolic programs. CLEVR-IEP [16] suggested predicting symbolic programs to be executed on a disentangled scene representation, followed by NS-VQA [17] and NS-CL [18]. V2A [19] introduced a symbolic approach for robotic tasks using only the initial state of the scene and was evaluated in a real robotic environment [6], however, only open-loop single-step tasks were demonstrated. Symbolic approaches were also considered in Task and Motion Planning systems (TAMP) that were often based on the formal language PDDL [20]. Classic TAMP approaches rely on predefined rules and known dynamic models [21, 22, 23], but learning-based methods are increasingly replacing handcrafted heuristics in TAMP methods [24, 25, 26]. PDDLStream [27] extends PDDL to add any sub-symbolic model in a black-box way, which also opens it to the use of learning-based methods. In [5], it was used for online replanning of short-horizon robotic tasks in the real world. Deep Visual Reasoning [28] directly predicts task plans by considering initial observation only. Regression Planning Networks (RPN) [29] propose a task planning model performing regression in symbolic space using neural networks. RPN are expanded in [30] by utilising regression planning on scene graph representations to estimate the next subgoal based on the current scene and the scene graph, i.e. solving only a subpart of the overall task that our method deals with. The interaction with the real world that follows a linguistic query (e.g.measure the weight of the mug; find the sofa) is explored in embodied question answering approaches [31, 6, 32]. In [6], language-conditioned visual reasoning is combined with manipulation. Navigation of agents based on vision and language was investigated in VLN [1] or [33]. However, both approaches considered tasks that do not react to incoming measurements, i.e. reasoning not conditioned on new observations. Recently, several works utilised large-language models (LLMs) to generate action plans for robot manipulation tasks (e.g. [34], [35], [36], [12], [37]). While they show impressive versatility and adaptability to novel environments, they suffer from a lack of fast-loop responsiveness, verifiability, and repeatability. LLMs require careful prompt engineering and filtering of outputs before the execution (e.g. [11]). Furthermore, such approaches do not capture the task and trajectory details and do not enable fast-loop reactiveness. These problems have to be resolved before broader LLMs deployment. Using data created with MuBlE environment, we train and evaluate a novel CLIER reasoning agent that can act on a natural language query in a real scene. Moreover, we present closed-loop reasoning experiments in simulation and real tabletop scenes. A highlight of our approach is that it plans for the perception of invisible object properties, e.g. when stacking objects by weight. Moreover, unlike classical methods, our reasoning agent is executed on every keyframe, which allows it to recover from a wide range of manipulation failures (e.g., unsuccessful grasping or an object falling during manipulation).
III The Proposed CLIER Method
@@
In this section, we present our closed-loop interactive reasoning approach (CLIER) for long-horizon manipulation tasks that require embodied perception. The main modules are presented in Fig. 1 and an example task in Fig. 4. CLIER method is implemented within the MuBlE simulation environment [14], which is introduced in Sec. IV-A.
In the following we introduce the individual parts of our reasoning pipeline. CLIER starts by parsing the visual scene (see Sec. III-A) to create or update its semantic and geometric scene graph (see Sec. III-B). As the reasoning agent we propose a multi-staged approach inspired by Neural Symbolic VQA approaches [38, 17]. Simultaneously to the scene parsing, a symbolic program is extracted from a natural language instruction (see Sec. III-C). Afterwards, the Symbolic Program Executor evaluates the symbolic program on the scene graph to select the subgoals that have to be achieved on the scene, e.g. measuring the weight of the target when the graph node of the target’s weight is empty (see Sec. III-D). Such a subgoal is forwarded to the Action Planner, which takes into account the current state of the scene graph and the given subgoal, and selects a sequence of primitive actions to be executed in the environment. Each primitive action is executed within the physics loop (see Sec. III-E). After each primitive action, the Action Planner reevaluates the next step within the action (render) loop (see Sect. III-F). The reasoning is completed when the Symbolic Program Executor finishes evaluation and either the answer is obtained, or the state of the scene matches the target state in the task completion modules.
@@
Figure 3: (Left) Examples of simulated visual observations (selected frames) generated for the actions corresponding to the instruction: Stack the lightest of metal objects on the yellow object. Note that metal cans were initially picked up to measure their weight. Further, according to the measurement taken, the heavier of the cans was put down and the lighter one was picked up and stacked on the yellow plate. (Right) Example simulated scenes and corresponding instructions in natural language generated with MuBlE (in the dataset, instructions left to right belong to tasks 7, 3, and 1 in Tab. I).
III-A Scene Parser
@@
A real scene can be captured by a camera, or a synthetic scene can be rendered by Blender based on a detailed scene description within the MuBlE environment (see Sec. IV-A). We capture/render the scene in every keyframe - i.e. after a motion corresponding to the primitive action is finished (e.g. moving from object 1 to object 2, approaching grasping pose, closing the gripper, etc.). The keyframe-based approach allows us to use Blender for rendering frames, which have sufficient quality for sim2real transfer, and sufficient speed to allow for recovering from failures with the closed-loop approach. Example key frames rendered during execution of a task are presented in Fig. 3 (Left).
@@
To parse the scene into a set of objects with attributes $\mathtt{\tilde{G}}$, we apply segmentation with Mask R-CNN [39], followed by classification of object attributes with ResNet [40] as in SHOP-VRB [38] and pose regression with CosyPose [41].
III-B Scene Graph
A Scene graph $\mathtt{G}$ is generated from the parsed scene for each key frame $\mathtt{I}$ starting from the first timestep (see Fig. 4). The Scene graph consists of feature vectors corresponding to each object in the scene, including categorical attributes for every object (e.g. material, colour), along with their positions and orientations and physical properties (e.g. weight, stiffness, elasticity, etc.). It also includes information on whether the object is in the gripper or whether the gripper is closed on it. Please note that the value of some of these properties might be empty (e.g. unknown weight of the object).
We use geometrical heuristics to create or update the scene graph $\mathtt{G}$ for keyframe $\mathtt{I}$, similarly to [30]. These heuristics compare the relative positions of the objects and the end effector in order to define the following attributes: whether the object is in the gripper, whether the object is within the feasible grasp, whether the object is raised above the table, and whether the gripper is placed directly above the object.
III-C Symbolic Program Generator
The first stage of our proposed reasoning approach inspired by neuro-symbolic VQA approaches [38, 17] is a generation of a symbolic program. Symbolic Program Generator is implemented as a Seq2Seq network [42] translating the natural language query $\mathtt{T}$ into a sequence of symbolic programs similarly to [38], trained with ground-truth. The generated program is in the CLEVR-IEP [16] format.
Given the program sequence, it is evaluated on the current state of the scene graph $\mathtt{G}$, and either the task is finished, or a subgoal $\mathtt{S}$ necessary for task completion is produced, e.g. measuring the weight of the target when the graph node of the target’s weight is empty. When the Action Planner achieves the subgoal, the program is reevaluated and produces the next subgoal until task completion.
III-D Action planner
The second stage of the reasoning is an Action Planner that for every keyframe $\mathtt{I}$ and a subgoal $\mathtt{S}$, given the scene graph $\mathtt{G}$, produces a sequence of primitive actions $\mathtt{[A]}$.
We employ a 2-layer transformer encoder as the action planner to predict the next primitive action $\mathtt{A}$ and its target, given the current scene graph $\mathtt{G}$ and the sub-goal $\mathtt{S}$, e.g. measuring the weight of the target when the graph node of the target’s weight is empty. We construct the input as the concatenation of the embeddings of the goal token and vectors describing all objects in the scene (including categorical properties, pose information, and relation to gripper based on the aforementioned heuristics). From this encoding, we predict the next action token, and select the target using the attention distribution over the objects training the network with ground-truth actions supervision. For example, when the symbolic program is queried for the task goal filter_weight[heaviest], it encounters an empty field weight in one of the scene graph nodes. This triggers the corresponding subgoal $\mathtt{S}$ for the Action Planner (i.e. measure_weight with a pointer to the graph node of the object with the empty field).
Further, action planning regresses the task to a sequence of primitive actions for the current state of the scene. The Action Planner takes the current scene graph, the subgoal and its target as the input and produces the sequence of primitive actions that lead to executing the subgoal—e.g. $\mathtt{move}$$\mathtt{approach}$$\mathtt{close\_gripper}$$\mathtt{lift}$$\mathtt{weigh}$ for measure_weight subgoal). These actions are executed within the physics loop (see Sec. III-E). Non-visual states, such as weight or stiffness are encoded in the scene graph after measurements, however, the transformer is provided only a binary flag indicating whether the property was measured or not and the reasoning is performed once all the measurements are completed. The binary flags simplify the process as the reasoning can be done simply over the binary values instead of continuous measurement outputs.
III-E Primitive actions and physics loop